OpenStudy (anonymous):

A gas having molecular formula On.if its vapour density is 24.the value of n is

OpenStudy (mos1635):

need more input..... O stunds for oxygen 24. in what?

OpenStudy (anonymous):

mol. wt . = 2 * V.D. mol wt=48 On will be o3 as 16*3 = 48